OpenStudy (anonymous):

Add: (4 – 3n – 5n2) + (–2 + n – 3n2) A. 2 – 4n – 8n2 B. 2 – 2n – 8n2 C. –2 – 2n – 8n2 D. –2 – 4n – 8n2 I believe this one is B.

OpenStudy (anonymous):

Add: (2x – y – 2z – 6) + (–6x + 4y + 3z + 8) A. –4x + 3y + z + 2 B. 4x – 3y – z + 2 C. –4x + 3y –z – 2 D. 4x – 3y + z + 2 I believe this one is A.

OpenStudy (mathstudent55):

Let's do the first problem first.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

So for number 1 here are the steps I did...

OpenStudy (anonymous):

4 + -2 = 2 -3n + n = -2n -5n^2 - 3n^2 = -8n^2

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

And then I put them all together so equal 2 - 2n - 8n^2

OpenStudy (anonymous):

Okay great!

OpenStudy (mathstudent55):

I prefer to keep rewriting what the problem is, so I don't forget any terms. This is how I'd do it.

OpenStudy (anonymous):

That makes sense. Is #2 correct?

OpenStudy (anonymous):

@imqwerty

OpenStudy (mathstudent55):

\((4 - 3n - 5n2) + (-2 + n-3n^2) \) Since the two trinomials are being added, the parentheses can be dropped. \(= 4 - 3n - 5n^2 + (–2) + n - 3n^2 \) Now we can place like terms together. \(= 4 + (-2) - 3n + n - 5n^2 - 3n^2 \) \(= 2 - 2n - 8n^2\)

OpenStudy (mathstudent55):

Ok. Let's look at 2.

OpenStudy (anonymous):

Oh that's a smart way of doing it. I never thought about that.

OpenStudy (mathstudent55):

\( (2x - y - 2z - 6) + (-6x + 4y + 3z + 8)\) Same idea here. We are adding polynomials, so we can drop both sets of parentheses. \(= 2x - y - 2z - 6 + (-6x) + 4y + 3z + 8\)

OpenStudy (mathstudent55):

Now we move the like terms together: \(= 2x + (-6x) - y + 4y - 2z + 3z - 6 + 8\) Now we combine each set of like terms: \(=-4x + 3y + z + 2\)

OpenStudy (mathstudent55):

You are correct again.

OpenStudy (anonymous):

Oh great. Thanks for helping. I only have a few more left if you don't mind. You are a lifesavor!